# MinD-3D: Reconstruct High-quality 3D objects in Human Brain

### v1

@@ Appendix A More Implementation Details
A.1 Neuro-Fusion Encoder
Our encoder consists of the encoder from the pre-trained LEA [26] and a feature aggregation module. LEA, a transformer-based auto-encoder, encodes fMRI signals and reconstructs them as depicted in Fig. A.1. Unlike the original Masked Autoencoder (MAE) [16], which only masks some patch-based latent codes, LEA masks all such codes. This enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token during training, thereby facilitating the development of a dense fMRI representation within the [CLS] token.
LEA is trained on the UKB dataset [23]. As shown in Fig. A.1, when LEA is directly applied to our fMRI-Shape dataset, it yields rough reconstructions. However, these reconstructions are not highly precise, primarily due to variations in scanners, subjects, and tasks.
Figure A.1: Architecture of LEA. During training, LEA masks all the patch-based latent codes and enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token.
Given that our input fMRI signals are in video form, to efficiently extract both semantic and structural features from this complex data, we opt to fine-tune the encoder with a small learning rate during the initial stage of training with hyperparameters provided in Tab. A.1.
@@ A.1 Neuro-Fusion Encoder
Table A.1: Hyperparameters used in neuro-fusion encoder.
A.2 Latent Adapted Decoder
Our decoder is adapted from Argus [39]. The original Argus model enhances the quality and diversity of generated 3D shapes by incorporating tri-plane features as latent representations and a discrete codebook for efficient quantization, leveraging the power of transformers for multi-modal conditional generation. It generates latent codes in an auto-regressive manner, similar to VQGAN [8], based on conditional features. Subsequently, it decodes the quantized vectors into point clouds.
In our latent adapted decoder, we add an adapter layer after every 8 transformer blocks to fuse the fMRI feature. During the second stage of training, we focus on optimizing the parameters in the adapter layers and the learnable parameter $\lambda$. The hyperparameters of our latent adapted decoder are detailed in Tab. A.2.
@@ A.2 Latent Adapted Decoder
mlp ratio 0.99 sequence length 1024
Table A.2: Hyperparameters used in latent adapted decoder.
Appendix B FMRI-Shape Details
To better illustrate our proposed dataset, fMRI-Shape, we provide a detailed description in this section.
Figure B.1: Examples of 360-Degree View Videos. These videos effectively showcase the intricate details, lighting, and shadows of the 3D objects. Presented in a sequence from left to right, they depict the order in which the subjects viewed them. Each video boasts a high resolution of 1024×1024, providing a clear and detailed visual experience.
B.1 More Experimental Details
To clarify the experimental details, we showcase our specially rendered 360-degree view videos. In Fig. B.1, six frames are selected from four videos. The camera in each video encircles the object, completing a full 360-degree clockwise rotation, forming an 8-second video with 192 frames. These high-resolution videos (1024×1024 pixels) vividly capture the intricate details, lighting effects, and shadows of the 3D objects. The arrangement of the videos mirrors the order in which they were viewed by the subjects. Displayed on an LCD screen, these videos enable the subjects to observe and comprehend all facets of the 3D objects. This viewing strategy is crucial, as it aids the subjects’ brains in accurately perceiving the objects, a fundamental aspect for gathering precise and high-quality fMRI signals that correspond to the visual stimuli observed by the subjects.
B.2 More Preprocessing Details
In preparing our data for training, we provide a visualization that clearly shows how we identify the visual Regions of Interest (ROIs) in our data. The preprocessing is applied to all fMRI frames.
Figure B.2: Selection Process of Visual ROIs: On the left, the fMRI is depicted in its original signal form, while on the right, the corresponding area is transformed into a 2D image, demonstrating the conversion process.
Refer to Fig. B.2. The first row displays the raw fMRI signals and their 2D image formation. In the second row, on the left, you can see the appearance of the fMRI signal and the locations of the ROIs. On the right, there is a 2D image illustrating where these ROIs are located within the signal, which has been converted from a format known as 32k_fs_LR.
After identifying these ROIs in the 1023×2514 2D fMRI data, we resize them to 256×256, as shown in Fig. B.3. It’s important to note that our Mind-3D model uses these resized, single-channel 2D fMRI signals as its input.
Figure B.3: Converting ROIs in fMRI Signal
Appendix C More Results
Subject Semantic-Level Structure-Level
[…]
sub-001 0.842 0.424 0.411 0.421 0.704 0.212 0.478 0.069
sub-002 0.836 0.424 0.412 0.420 0.705 0.218 0.483 0.062
sub-003 0.829 0.427 0.410 0.416 0.703 0.212 0.479 0.064
Table C.1: Quantitative results for all subjects.
Qualitative results from different views. In Fig. C.1, we showcase a collection of 2D rendered images of 3D objects by MinD-3D from various viewpoints for subject-1 in fMRI-Shape, closely resembling the ground truth (GT) objects both in semantics and structure.
Quantitative results for all subjects. Our proposed dataset, fmri-shape, includes data from three subjects. To demonstrate the robustness of our model, we present comprehensive quantification results for all three subjects. Tab. C.1 showcases these full metrics, indicating that our model consistently performs well across different subjects. These results underscore the model’s robustness in handling subject variability.
Figure C.1: Qualitative results for subject-1 from various viewpoints.
Appendix D Limitations and Future Works
As the first attempt at the task of Reconstructing 3D Objects from Mind in fMRI (Recon3DMind), our MinD-3D has successfully reconstructed 3D objects that are semantically and structurally similar to the originals. However, the reconstructed 3d objects lack detailed texture, color, and other aspects compared to the original point clouds. To overcome this limitation, our future work will focus on developing methods to more effectively extract texture and color details from 3D objects within fMRI signals. We aim to refine our reconstruction method to create more accurate and detailed replications of the original 3D objects, thereby enhancing the fidelity of the reconstructed models and capturing the complexity and richness of the original designs. Regarding our dataset, we recognize that relying solely on the fMRI-Shape dataset is insufficient for the task’s full scope. A notable challenge is mitigating dataset noise due to the cognitive biases individuals might hold towards certain objects. Therefore, it’s necessary for us to expand our fMRI-Shape dataset by including more 3D objects and collecting data from additional subjects, with the hope of advancing research in our proposed Recon3DMind field.

### v3

@@ Appendix 0.C More Implementation Details
0.C.1 Neuro-Fusion Encoder
Figure 0.C.1: Architecture of LEA. During training, LEA masks all the patch-based latent codes and enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token.
Our encoder consists of the encoder from the pre-trained LEA [34] and a feature aggregation module. LEA, a transformer-based auto-encoder, encodes fMRI signals and reconstructs them as depicted in Fig. 0.C.1. Unlike the original Masked Autoencoder (MAE), which only masks some patch-based latent codes, LEA masks all such codes. This enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token during training, thereby facilitating the development of a dense fMRI representation within the [CLS] token.
LEA is trained on the UKB dataset [30]. As shown in Fig. 0.C.1, when LEA is directly applied to our fMRI-Shape dataset, it yields rough reconstructions. However, these reconstructions are not highly precise, primarily due to variations in scanners, subjects, and tasks.
Given that our input fMRI signals are in video form, to efficiently extract both semantic and structural features from this complex data, we opt to fine-tune the encoder with a small learning rate during the initial stage of training with hyperparameters provided in Tab. 0.C.1.
Table 0.C.1: Hyperparameters used in neuro-fusion encoder.
@@ 0.C.1 Neuro-Fusion Encoder
0.C.2 Feature Bridge Diffusion Model
We use a transformer-based diffusion model to generate vision features. During the inference phase of the diffusion model, we concatenate the fMRI token, the time token, the predicted image token, and the learnable token. The output learnable token then serves as the predicted image token for the model’s next inference. The hyperparameters of our feature bridge diffusion model are shown in Tab. 0.C.2.
Table 0.C.2: Hyperparameters used in feature bridge diffusion model.
parameter value parameter value
depth 6 heads 8
embedding dim 512 timestamps 100
mlp ratio 0.99 token length 770
0.C.3 Latent Adapted Decoder
Our decoder is adapted from Argus. The original Argus model enhances the quality and diversity of generated 3D shapes by incorporating tri-plane features as latent representations and a discrete codebook for efficient quantization, leveraging the power of transformers for multi-modal conditional generation. It generates latent codes in an auto-regressive manner, similar to VQGAN, based on conditional features. Subsequently, it decodes the quantized vectors into 3D mesh.
In our latent-adapted decoder, we add an adapter layer after every two (T=2) Transformer blocks to fuse the fMRI feature. During the second stage of training, we freeze the parameters of Argus and focus on optimizing the parameters in the adapter layers. The hyperparameters of our latent adapted decoder are detailed in Tab. 0.C.3.
Table 0.C.3: Hyperparameters used in latent adapted decoder.
@@ 0.C.3 Latent Adapted Decoder
mlp ratio 0.99 sequence length 1027
Appendix 0.D Features Analysis
To understand the biological relevance of our feature extractor, in line with previous studies [51, 47, 32], we employ a linear encoding model to project the extracted fMRI features $c_{f}$ onto brain activity patterns, using subject 1 as a representative case for analysis.
Figure 0.D.1: Quality of the extracted features. We map the features to brain regions via ridge regression modeling using $c_{f}$, $c_{v}$ and $\hat{c_{v}}$, with voxel-wise Pearson’s correlation coefficients computed between them. The visual function ROI is highlighted in red. In contrast, other regions are depicted in blue.
Specifically, we utilize ridge regression to establish the mapping between $c_{f}$ and the fMRI signals within the training dataset. We then project $c_{f}$ from the test set into the fMRI space and compute the Pearson correlation coefficient to assess the relationship. As illustrated in Fig. 0.D.1, areas positively correlated with visual Regions of Interest (ROIs) are marked in red, while other regions showing positive correlations are highlighted in blue. The results demonstrate a strong positive correlation between $c_{f}$ and visual ROIs, confirming the effective extraction of relevant features. Additionally, we perform a similar correlation analysis with the visual latent feature $c_{v}$, which also shows positive correlations. Notably, the correlation strength of $c_{f}$ surpasses that of $c_{v}$, indicating its superior relevance to brain activity. However, it is important to note that in some measures, the correlation coefficient of $c_{v}$ is higher than that of $c_{f}$, which highlights the complexity of the task and underscores the critical role of bridging features. Furthermore, our results indicate that our generated visual latent features achieve activation levels comparable to the ground truth (GT), emphasizing the effectiveness of our comparative learning approach and demonstrating well-aligned fMRI feature and image spaces.
Appendix 0.E Future Works
As the first attempt at the task of Reconstructing 3D Objects from the Mind using fMRI (Recon3DMind). Our proposed MinD-3D model has successfully reconstructed 3D objects that are semantically and structurally similar to their originals, thereby proving the feasibility of the task. However, our model’s 3D representations, which are mesh-based, accurately outline the shape of objects but do not capture color and other details as precisely as the original objects. To improve on this, we plan to develop new methods for more effectively extracting texture and color information from fMRI signals. Our goal is to refine our reconstruction technique to produce more detailed and accurate reproductions of the original 3D objects, thereby improving the quality of the models and more faithfully capturing the originals’ complexity and nuances.
Crucially, we introduce the first 3D vision fMRI dataset with contributions from 14 participants across various settings. Currently, most fMRI-based visual reconstruction methods are specific to individuals, making out-of-distribution (OOD) testing, both across individuals and classes, extremely challenging. Our proposed fMRI-Shape dataset presents an opportunity to tackle this formidable challenge. Addressing individual differences is a complex task that necessitates a collaborative effort from the research community with the goal of enhancing the adaptability of such models. Regarding our dataset, we acknowledge that relying solely on the fMRI-Shape dataset is not sufficient for comprehensively addressing the task at hand. Therefore, we plan to enrich the dataset by sampling more 3D data, including all objects from ShapeNet, and involving more participants. This effort aims to expand the fMRI-Shape dataset and further advance research in fMRI-based human brain decoding.